OpenStudy (anonymous):

What is the square root of 50 times the square root of 45?

OpenStudy (anonymous):

47.4341649025

OpenStudy (anonymous):

\[\sqrt{50}*\sqrt{45}=\sqrt{50*45}=\sqrt{5*2*5*5*3*3}=\sqrt{5*2}=\sqrt{10}\]

OpenStudy (cgreenwade2000):

^No. Tom.

OpenStudy (anonymous):

\[5*3*\sqrt{10}=15\sqrt{10}\] forget the stuff in front of the root

OpenStudy (anonymous):

got a little to excited XD